The following data reflect the current month’s activity for Vickers Corporation.

Actual total direct labor $ 639,115
Actual hours worked 36,500
Standard labor-hours allowed for actual output (flexible budget) 35,300
Direct labor price variance $ 17,885 F
Actual variable overhead $ 159,400
Standard variable overhead rate per standard direct labor-hour $ 4.40

Variable overhead is applied based on standard direct labor-hours allowed.

Required:

Compute the labor and variable overhead price and efficiency variances. (Indicate the effect of each variance by selecting "F" for favorable, or "U" for unfavorable. If there is no effect, do not select either option.)

Direct labor: Price variance Efficiency variance Variable overhead: Price variance Efficiency variance
Answer #1

Direct labor price variance = (standard rate - actual rate)*actual hours

17885 = (standard rate - 639115/36500)*36500

Standard rate = $18 per hour

Direct labor efficiency variance = (standard hours - actual hours)*standard rate

= (35300-36500)*18

= $21600 unfavorable

Variable overhead price variance = (4.40 - 159400/36500)*36500 = $1200 favorable

Variable overhead efficiency variance = (35300-36500)*4.40 = $5280 unfavorable